OpenStudy (anonymous):

multiply and simplify (√5-√8)(√5+√8) (√5)^2(√8)^2 25-16 = 9 please help is this correct?

OpenStudy (anonymous):

oops 25-64=39 simplify???

OpenStudy (jamesj):

(√5-√8)(√5+√8) = √25 - √64 = 5 - 8 = -3

OpenStudy (anonymous):

Oh okay thanks so much!!!